### v1

# A Non-Abelian Gauge Theory for Surface Excitations of ${}^{3}$e-B

Abstract
In ${}^{3}$e-B, two atoms pair in an orbital angular momentum $1$ spin triplet state above the phase transition temperature with $SO(3)\times SO(3)$ symmetry. Below the transition temperature, this symmetry is spontaneously broken to the diagonal $SO(3)$ due to spin-orbit coupling. Considerations based on effective potentials and solitons show that $SO(3)$’s gets enhanced to $SU(3)$’s and the symmetry breaking is that of $G=SU(3)\times SU(3)\times U(1)$ to $H=SU(3)$. The theory of the resultant Goldtsone modes can be naturally formulated as a gauge theory of $H$. Its Gauss law is treated here and shown to lead to surface states in a container with a dynamics governed by large gauge transformations.
Observable consequences are pointed. The transference of the analysis to the chiral model of QCD is pointed out where $SU(3)$ are the left and right chiral groups, $U(1)$ is the axial $U(1)$ , the surviving symmetry is flavour $SU(3)$ and the Goldstone modes are the pions.
1 Introduction
In ${}^{3}$e-B , the superfluid transition takes place by a pairing of this molecule in its $P$-wave spin triplet state with another such molecule in an identical state.
The order parameter for this transition can be denoted by a complex $3\times 3$ matrix of fields $\phi=\{\phi_{s}(x)\}$ following Mermin [1] where $SO(3)$ of spin acts on $\phi$ by left multiplication and $SO(3)$ of orbital momentum acts on $\phi$ right multiplication.
The chosen dynamics for $\phi$ is invariant under these actions of $SO(3)_{L}\times SO(3)_{R}$, $L$ and $R$ denoting the left and right groups.
But this symmetry is spontaneusly broken to the diagonal $SO(3)_{J}$ of total angular momentum by the boundary condition
$\lim_{|\vec{x}|\rightarrow\infty}\phi(x)=\mathbb{I}$ (1.1)
Thus there is a spontaneous breakdown of the group $G=SO(3)_{L}\times SO(3)_{R}$ to the diagonal $H=SO(3)_{J}$. This breakdown is induced by the $L-S$ coupling.
This symmetry breaking pattern is reminiscent of the chiral symmetry group $SU(3)_{L}\times SU(3)_{R}$ of massless $QCD$ with three flavours and its spontaneous breakdown to the diagonal $SU(3)_{L+R}$. We can in fact bring this resemblance closer by noting that $G$-invarint potentials are traces of polynomials in $\phi^{\dagger}\phi$ of the form
$\Sigma_{n}a_{n}\mathrm{Tr}(\phi^{\dagger}\phi)^{n}.$ (1.2)
These potentials are invariant if $SO(3)_{L,R}$ is enlarged to $\mathcal{G}=SU(3)_{L}\times SU(3)_{R}\times U(1)$ where $U(1)$ acts by an overall phase: $\phi\rightarrow\omega\phi,|\omega|=1$. Then the vacuum symmetry group is $H=SU(3)_{L+R}$. These are our groups of interest.
The $U(1)$ group above is the analogue of the axial $U(1)$ of $QCD$. Just as in $QCD$, it is spontaneously broken here as well.
$G$ does not act effectly on $\phi$. If $c=e^{\frac{2\pi i}{3}\mathbb{I}_{3}}$, with $\mathbb{I}$ being the $3\times 3$ identity matrix of $SU(3)$, generates the center of $SU(3)$, then the order $9$ discrete group $Z_{9}$ with elements diag$\{c^{m},\bar{c}^{n},\bar{c}^{n-m}\}$ with $n,m$ belonging to the set $\{1,2,3\}$ acts trivially on $\phi$. Thus in quantum theory only representations of $\mathcal{G}$ where $Z_{9}$ becomes identity are allowed. With this proviso, we can work with $G$.
In Mermin’s review , $\phi$ is restricted to $\omega\times R$ where $R$ is real and orthogonal. But a previous work [2] has found that for discussions with solitons of ${}^{3}$e-B it is important to consider $\phi$ to be any $3\times 3$ matrix ( with the boundary condition mentioned above). This paper will proceed in this latter direction.
2 Gauge Theory of Goldstone Modes
The next step in this paper is to describe the symmetry breakdown $G\rightarrow H$ as a gauge theory for $H$ for fields valued in $G$. This is an old result [3] reviewed below. What will be new here will involve the description of its superselection sectors following [4], [5], [6] in sec. 3. This discussion involves new treatment of Gauss law and observables described in [4].
We will see that for finite containers, they induce surface states and their dynamics as in “topological insulators”. Just as in the latter, their origin is from spin-orbit coupling.
Is ${}^{3}$e-B a model in the lab for flavour physics of QCD?
Maybe so. It appears to have all the features of asymptotic QCD. But QCD has coloured fields like quarks.
What are their analogues here?
3 Spontaneous Breakdown as a Gauge Theory
The paper of Balachandran, Stern and Trahern [3] which treats this approach is summarised here. The discussion of the quantum generators of this gauge group, and its superselection sectors are reserved for the next section.
The gauge groups $\mathcal{G},\mathcal{H}$ are the maps from spacetime to Lie groups $G,H$ in our considerations. , they are the restrictions of $\mathcal{G},\mathcal{H}$ to constant time, say 00 , that will be relevant. They are denoted by the same symbols. We consider four dimensional $M$ so that the constant time slices $M_{3}$, are three dimensional. It can be $\mathbb{R}^{3}$ or the 3-ball $B_{3}$ with coordinates $x$ and the boundary $S^{2}$. The evaluation of their elements at $x$ will be denoted by $g(x)$ and $h(x)$, respectively. Often we will suppress $x$.
The adaptations of this paper’s arguments to any dimension is straightforward.
Let $\underline{G}$ and $\underline{H}$ denote the Lie algebras of $G$ and $H$ with basis {$L_{\lambda}$} and {$T_{\alpha}$}, in the defining representation of $G$ with $\mathrm{Tr}(L_{\lambda}L_{\lambda^{\prime}})=2\delta_{\lambda\lambda^{\prime}}$,Tr($\mathrm{Tr}(T_{\alpha}T_{\alpha^{\prime}})=2\delta_{\alpha\alpha^{\prime}}$.
The adjoint action of $H$ leaves $\underline{H}$ invariant:
$hT_{\alpha}h^{-1}=h_{\beta,\alpha}T_{\beta}$ (3.1)
where the matrix of this transformation in the basis {$T(\alpha)$} of $\underline{H}$ has also been denoted by $h$.
Hence it also leaves its orthogonal complement $\underline{G/H}$ in $\underline{G}$ invariant . If {$S(i)$} is its basis with
$\{S(i):\mathrm{tr}~{}S(i)S(j)=2\delta_{ij},$ (3.2)
we can write
$hS(i)h^{-1}=[D(h)]_{ji}S(j)$ (3.3)
where $D(h)$ is real and orthogonal. We sum on repeated indices here and below.
If $v$ is in $\underline{H}$,then
$v=\frac{1}{2}T(\alpha)\mathrm{Tr}(vT(\alpha)),$ (3.4)
while if $w$ is in its orthogonal complement ,
$w=\frac{1}{2}S(j)\mathrm{Tr}(wS(j)).$ (3.5)
The above equations are clearly valid if $h\in H$ is replaced by $h(x)$.
Now consider
$A_{\mu}(g(x))=\frac{i}{2}T(\alpha)\mathrm{~{}Tr~{}}T(\alpha)g^{-1}(x)\partial_{\mu}g(x)$ (3.6)
Under the transformation
$g(x)\rightarrow g(x)h(x)$ (3.7)
$A_{\mu}$ transforms as a connection for the gauge group $\mathcal{H}$:
$A_{\mu}(gh)=h^{-1}A_{\mu}h+ih^{-1}\partial_{\mu}h.$
This follows on using the orthogonality of $Ad~{}H$ and the identity (3.1) with $v_{\mu}=h^{-1}\partial_{\mu}h$ ( $x$ has been suppressed).
In contrast
$B_{\mu}(g)=\frac{i}{2}S(j)\mathrm{~{}Tr~{}}S(j)g^{-1}(x)\partial_{\mu}g(x)$ (3.8)
transforms tensorially despite the derivative on $g$:
$B_{\mu}(gh)=h^{-1}(x)B_{\mu}(g)h(x)$ (3.9)
We can now consider $\mathcal{H}$ invariant interactions like
$\mathrm{tr}F~{}^{*}F,\qquad\mathrm{Tr~{}}B_{\mu}B^{\mu}$ (3.10)
etc for the dynamics of the Goldstone modes, where $F$ is the field strength associated with the gauge connections $A$. But there are more invariants available for $\mathcal{G,H}$ of ${}^{3}$e-B . And there remains the treatment of Gauss law which is in the next section.
The gauge group $\mathcal{H}$ acts on $g_{L,R}$ according to $g_{L,R}\rightarrow g_{L,R}h$. Hence,
$g_{L}g^{\dagger}_{R}\equiv\phi=\left\{\phi_{s,i}\right\}$ (3.11)
is gauge invariant while $s$ and $i$ carry $SU(3)_{L,R}$ actions. This is our order parameter.
Also gauge connections can be constructed for $\mathcal{G}_{L,R}$ separately giving two tensorial fields $B_{\mu~{}L,R}$ and gauge connections $A_{L},A_{R}$. So many sorts of interactions are possible.
4 Gauss Law and Superselection Rules
Let $A$ be a chosen connection. Following $QCD$, consider the action
$S=-\frac{1}{4g^{2}}\mathrm{Tr}F_{\mu\nu}F^{\mu\nu}+q\mathrm{Tr}A^{\mu}J_{\mu}$ (4.1)
where $F_{\mu\nu}$ is the curvature of $A$ and $J^{\mu}$ is the covariantly conserved matter current. The constants are as in $QCD$.
In view of ( 3.9) , $F$ transforms as $h^{-1}Fh$ under an $\mathcal{H}$ gauge transformation so that $S$ is invariant even if $h$ is time dependent.
So the generator of this gauge transformation must come from the Gauss law constraint.
The derivation of this constraint by quantising $g$ in the action is awkward. But we can get it simply by varying $g$ with an infinitesimal time dependent gauge transformation:
$\delta g=g\epsilon,~{}~{}\epsilon~{}\textrm{time dependent}$ (4.2)
where
$\epsilon=\epsilon^{\alpha}(x)\cdot T(\alpha),$ (4.3)
$\epsilon^{\alpha}$ being generic imaginary functions.
Only $A_{0}$ varies with $\partial_{0}\epsilon$ and that variations is
$\delta A_{0}=-(\partial_{0}\epsilon^{\alpha})T(\alpha).$ (4.4)
The corresponding variation in the Lagrangian is
$\frac{1}{2g^{2}}\int d^{3}x~{}tr(D_{i}F_{i0})(\partial_{0}\epsilon^{\alpha})T(\alpha).$ (4.5)
Since $D_{i}F_{i0}$ is $\underline{H}$ valued, and $\partial_{0}\epsilon^{\alpha}$ are generic, we get the Gauss law in the usual form :
$D_{i}F_{i0}\approx 0.$ (4.6)
We can also add the matter coupling
$\int d^{3}xA^{\mu}J_{\mu}$ (4.7)
to the Lagrangian which will enhance (4.6) by the matter term $J_{0}$.
We must check that L.H.S of (4.6) generates gauge transformations. and that requires the commutators of $A^{i}$ and $F_{0i}$. We can derive them as follows.
The analogue of the $pdq$ term can be read off from (4.1) . It comes from $\frac{1}{2g^{2}}\mathrm{tr}F_{0i}\partial_{0}A_{i}$ from varying $A_{i}$ so that as in $QCD$, $A^{i}$ and $E_{i}=\frac{1}{2g^{2}}F_{0i}$ are conjugate.
Then as in $QCD$ , the generator of small gauge transformations is
$Q_{H}(X)=\int d^{3}x\mathrm{Tr}(D_{i}XE_{i}+XJ_{0})$ (4.8)
[…]
Note that $E_{i}$ and $A^{i}$ are not gauge invariant. They are not observables , but are elements of the field algebra.
The classical Gauss law (4.6) has the quantum version of ‘small ’ gauge tranformations with generators $Q_{H}(X)$ which fulfill a) and b):
a)
$Q_{H}(X)|~{}\cdot~{}\rangle=0$ (4.10)
[…]
b) If $K$ is an observable
$\left[Q_{H}(X),K\right]=0.$ (4.11)
Then, $X$ being compactly supported , we recover classical Gauss law by partial integration , while (4.11) is to ensure that observables preserve the domain defined by (4.10). ( There are technical issues with domains which remain to be answered.)
But one can also consider
$Q(\Xi)=\int d^{3}x\mathrm{Tr}(D_{i}\Xi E_{i}+\Xi J_{0})$ (4.12)
where $\Xi$ are smooth functions which may not vanish at infinity ( or the boundary). These generate ‘large gauge transformations’ . As shown in previous papers, $Q(\Xi)$ commute with local observables due to local commutation relations , but in non-abelian theories such as ${}^{3}$e $B$ or $QCD$, they may not mutually commute. This has specific consequences for surface states as we shall see.
Two important points should be observed:
  1. 1.
    $Q(\Xi)$ generate a Lie algebra:
    $\left[Q(\Xi),Q(\Xi^{\prime})\right]=iQ([\Xi,\Xi^{\prime}]).$ (4.13)
    But the rhs can be a small gauge transformation, so that the large ones by themselves do not generate a Lie algebra.
  2. 2.
    Adding a small gauge transformation $Q_{H}(X)$ to $Q(\Xi)$ is still a large gauge transformation $Q(\Xi+X)$ which has the same action on a quantum state in view of (4.10):
    $Q(\Xi+X)|\cdot\rangle=Q(X)|\cdot\rangle$ (4.14)
    Further since $[\Xi,X]$ is a compactly supported Lie algebra-function, $Q([\Xi,X])$ is a small gauge transformation.
    So let $\mathcal{H}$ be the Lie algebra of all gauge transformations, large and small, and $\mathcal{H}_{loc}$ the Lie algebra of small ones. Then the latter is an invaraint subalgebra of the former and the large algebra is the quotient $\mathcal{H}/\mathcal{H}_{loc}$ is the Lie algebra of large gauge transformations. Its elements are equivalence classes $a+\mathcal{A}_{loc}$ where $a\in\mathcal{H}$.
[…]
    If $\psi$ is a gauge-noninvariant field such as a ${}^{3}$e-B or a quark field in $QCD$, and elements of $\mathcal{H}$ multiplies it, they are no longer gauge transformations , failing to fulfill requisite algebraic relations.
    Superselection Rules
    Let $\mathcal{A}_{loc}$ denote the local algebra of gauge invariant observables. Its elements commute with all elements of the gauge algebra of $\mathcal{H}$. In a customary notation , $\mathcal{H}=\mathcal{A}_{loc}^{\prime}$. So in an irreducible representation of $\mathcal{A}_{loc}$, one can diagonalise a complete commuting set of operators (CCS) from $\mathcal{H}$ and the center of its enveloping algebra. For ${}^{3}$e-B, they are picked from $Q(\Xi)$ where $\Xi$ on the boundary $S^{2}$ are maps of $S^{2}$ to $H=SU(3)_{L+R}$.
    We can illustrate these remarks by choosing the $8$ functions $\mathbb{I}\lambda_{\alpha}$ for $\Xi$ where $\mathbb{I}$ is the constant function with value $1$ on $\mathbb{R}^{3}$ and $\lambda_{\alpha}$ are the Gell-Mann matrices. Then, $Q_{\alpha}=Q(\mathbb{I}\lambda_{\alpha})$ fulfill the $SU(3)$ Lie algebra relation:
    $[Q_{\alpha},Q_{\beta}]=if_{\alpha\beta\gamma}Q_{\gamma}$ (4.15)
    for fixed Casimirs $c_{2.3}$ and a Cartan subalgebra with basis $Q_{3,8}$ , we can choose its eigenstates with eigenvalues $\tilde{\lambda}_{3,8}$ and an $SU(2)$ Casimir with eigenvalue $I(I+1)$ commuting with these $Q$’s:
    $|\tilde{\lambda}_{3},I,\tilde{\lambda}_{8},c_{2},c_{3}\rangle=|0\rangle\otimes|\tilde{\lambda}_{3},I,\tilde{\lambda}_{8},c_{2},c_{3}\rangle^{\prime}$ (4.16)
    This vector is the analogue of the vacuum vector in quantum field theory. On RHS, $|0\rangle$ is the Fock vacuum and the primed vector carries just the $SU(3)$ representation.
    The GNS construction then gives a Hilbert space representation of $\mathcal{A}_{loc}$.
    Matter quantum numbers have been omitted in (4.16).
    Sticking to $SU(3)$, its Lie algebra has plenty of elements which do not commute with $Q_{3,8}\mathbb{I}$, for example $Q_{1}=Q(\mathbb{I}\lambda_{1})$ if $\lambda_{~{}I}\neq 0$. This will change the vector (4.16) to a new one, but preserving $c_{2,3}$ As no element of $\mathcal{A}_{loc}$ can change $\lambda_{3,8}\mathbb{I}$, this vector by GNS gives a new representation of $\mathcal{A}_{loc}$.
    The non-abelian symmetry $SU(3)_{L+R}$ is thus ‘spontaneously broken’ to $U(1)\times U(1)$ with generators $Q_{3,8}$ or to $U(2)$ if $I=0$ ( [4], [5], [6]).
    These operators changing the diagonalised operators are also called ‘anomalous’ . That is because they do not change the equations of motion ( which are sensitive to only local values of $\Xi$ so that $Q(\Xi)$ act on them like small gauge transformations ) and so are classical symmetries, but change the representation of the local observables.
    Chiral ‘symmetry’ of $QCD$ is an example.
    A Remark on Effective Lagrangians and Anomaly Matching
    Effective theories are subject to anomaly matching of ’t Hooft [7] so as to reproduce the anomalies of the parent theory.
    The analysis outlined above shows that it is best to match the superselection algebras in the parent and effective theories.
    Surface states in ${}^{3}$e-B
    Consider ${}^{3}$e $B$ in say a spherical container. ( Containers with other topologies such as toroidal can also be treated).
    The symplectic form which makes $A_{i},E_{j}$ conjugate is valid for the action $\int F\wedge^{*}F$ for potentials such as those of Mermin [1] so that $Q_{H}(X)$ vanish on the quantum vector states $|~{}\cdot~{}\rangle=0$. This is the quantum Gauss law for the classical statement
    $D_{i}E_{i}=0.$ (4.17)
    Even in $QED$, this use of test functions is needed to express the generator of charge $U(1)$ as pointed out earlier.
    Recalling that $E_{i}$ are operator valued distributions, such smearing gives us operators on a Hilbert space and is also suggested by Sobolev’s treatment of elliptic operators on manifolds with boundaries.
[…]
    The field equations $F(\psi)=0$ involve only local fields $\psi$ and are gauge invaraint:
    $F(\psi)=0\Rightarrow F([Q(\Xi),\psi])=0$ (4.18)
    As $\Xi$ can be an $X$, (4.18) implies full gauge invariance even though $\psi$ may not commute with $Q(\Xi)$ as is the case with connections.
    Such a situation is familiar from $QED$ where the ‘chemical potential’ $N$ generates $U(1)$ transformations.
    Adding to the $QED$ Hamiltonian $H_{0}$, we work with the new Hamiltonian $H=H_{0}+N$ which gives exactly the same equations of motion, but affects the partition functions.
    For such reasons , let us call $Q(\Xi)$ as chemical potentials. If $H_{0}$ is the ${}^{3}$e-B Hamiltonian we start with where the vector state is the ‘vacuum’
    $|0\rangle,\qquad Q(\Xi_{i})|0\rangle=0,$
    we can go to the sector with $H$ quantum numbers as in (4.16). This is the analogue of the excited states considered by Doplicher, Haag and Roberts [8]. Expectation values of local operators change on the folium of states (4.16) . An example is given below.
    Let $\mu$ be a constant with dimension of energy. We can then introduce a chemical potential in a new Hamiltonian
    $h=h_{0}+\mu Q(\Xi)$ (4.19)
    where $Q(\Xi)$ does not commute with the $Q_{3,8}$. Then the vector state (4.16) will change and so will the expectation of the generic surface operators $Q(\Xi^{\prime})$. Thus one has dynamics on the two sphere $S^{2}$.
    For example, suppose $Q(\Xi)=aQ_{1}+bQ_{5}$ with $a/b$ being irrational. The time $t=0$ inital state (4.16) changes with time and is not periodic since $\exp(i\mu Q(\Xi))$ has no period. One shows that it keeps evolving the diagonalised operators from one to another Cartan subalgebra [6] . ( One can prove that it never returns to the initial state of $t=0$. ).
    The spectrum of $h$ will keep changing with time. If $\mu$ is switched off at time $T$, there will be edge transitions to the ground state. May be they can be seen.
    The surface operators $Q(\Xi)$ emerge from spin-orbit coupling as seen in the literature on topological insulators. But why are they ‘topological’ ? Perhaps because $Q(\Xi)$ has no spacetime metric in it. But they do have a Lie algebra metric.
    Interpretation of the Chemical Potential
    In $QED$, the term $N$ comes from coupling the system to an external source supplying the charge. The interpretation of $\mu Q(\Xi)$ is much the same, it comes from the measuring apparatus coupled to the system.
    Expectation values of local observables change when the superselection sector is changed. This happens already in QED.
    Expectation value of charge density $:J_{0}:$ vanishes in the vacuum sector, but it cannot vanish if $N$ is not an integral value $q$ in the charge $q$ sector. In this case, we have to construct $J_{0}$ from charged fields like the electron field. The last is not gauge invariant. It is associated with the field algebra.
    There are similar gauge invariant fields in any local gauge theory such as that of ${}^{3}$ He $B$. The operators $\mathrm{tr}(E^{2}(x)+B^{2}(x)),\mathrm{tr}E\cdot B$ are examples. Combination of such terms integrated gives the Hamiltonian in the absence of chemical potential and its expectation value keeps changing in time when the chemical potential keeps changing the vector state. Knowing their expectation values at time $t=0$ when the experiment starts and the chemical potential $\mu(t)Q(\Xi)$ is switched on ( it is now time independent ) and at time $T$, when it is switched off. one infers the properties of the local dynamics.
    An Important Remark
    The chemical potential $Q(\Xi)$ is intrinsically present in the Hamiltonian emergent from the action
    $S=\int d^{3}x~{}\mathrm{tr~{}}(-\frac{1}{4e^{2}}F_{\wedge}*F+A^{\mu}J_{\mu}).$ (4.20)
    The exact Hamiltonian is
    $\frac{1}{4e^{2}}F_{\wedge}*F+Q(A_{0}).$ (4.21)
    The standard treatments set $A_{0}=0$, but one sees that it plays the role of $\mu Q(\Xi)$. Its asymptotic value seems determined by experimentl setup: Having fixed the vector state, its evolution is governed by the chemical potential $Q(A_{0})$.
    In $QED$, $A_{0}$ is the electrostatic potential . By fluctuating it, the total energy gets fluctuated and causes observable transitions.
    Likewise, the chemical potential $\mu Q(\Xi)$ should cause observable effects and help probing the system.
    5 Final Remarks
    This paper is restricted to considering surface effects. Still it suggests observable effects such as energy fluctuations and transitions due to them from emission of probably Goldstone modes . In addition, the approach predicts solitons [2]. It predicts as well theta vacua which are superselection sectors which can be constructed for ${}^{3}$e $B$ following [5] and which are spinorial. Such matters are being omitted for brevity.
    Superselection sectors like the charged vectors in $QED$ can be constructed using fields $\psi$ ( or the Cuntz algebra following [8]). They will be needed for detailed calculations.
    The dynamics on the surface are determined by $Q(\Xi)$. The functions $\Xi$ can be $\Xi^{\alpha}(\hat{x})\lambda_{\alpha}$ on the surface where $Q(\Xi)$ dominate the dynamics. If $\Xi^{\alpha}(\hat{x})$ are constants on $S^{2}$ these being the case briefly considered above, the surface dynamics may be soluble. They seem related to the quantum KAM dynamics treated in [9].
    In that connection, note that in the example above with irrational $a/b$ in $aQ_{1}+bQ_{5}$ , the evolution under conjugation cannot bring back $Q_{3,8}$ to the same Cartan subalgebra, as only automorphisms of the cartan subalgebra are the elements of the Weyl group, which is $S_{3}$ where the elements have periods $2,3$ or 6. So if $\exp(i\tau(aQ_{1}+bQ_{5}))$ maps back $Q_{3,8}$ to the same Cartan subalgebra, one of these periods has to be an integer multiple of $6\tau$. But that is not possible.
    The considerations of this paper are valid for the $QCD$ chiral model on $R^{3}$ ( or $S^{3}$) . They remain to be studied.
    6 Acknowledgements
    This paper could not have been written without the help of Arshad Momen. I thank him. Helpful comments from Arshad,l Parameswaran Nair,Al Stern and Rakesh Tibrawala are also appreciated. I also thank the Institute of Mathematical Sciences,Chennai and the hospitality of my colleagues there ( especially the Director Ravindran and S. Digal) where this work was completed.